A DPF ANALYSIS YIELDS QUANTUM MECHANICALLY ACCURATE ANALYTIC POTENTIAL EN-ERGY FUNCTIONS FOR THE A^{1}Σ^{+} and X^{1}Σ^{+} STATES OF NaH

\underline{ROBERT J. LE ROY}, SADRU WALJI, KATHERINE SENTJENS, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Alkali hydride diatomic molecules have long been the object of spectroscopic studies.^{a} However, their small reduced mass makes them species for which the conventional semiclassical-based methods of analysis tend to have the largest errors. To date, the only quantum-mechanically accurate direct-potential-fit (DPF) analysis for one of these molecules was the one for LiH reported by Coxon and Dickinson.^{b} The present paper extends this level of analysis to NaH, and reports a DPF analysis of all available spectroscopic data for the A^{1}Σ^{+} − X^{1}Σ^{+} system of NaH which yields analytic potential energy functions for these two states that account for those data (on average) to within the experimental uncertainties.

aW.C. Stwalley, W.T. Zemke and S.C. Yang, J. Phys. Chem. Ref. Data 20, 153-187 (1991).

b J.A. Coxon and C.S. Dickinson, J. Chem. Phys. 121, 8378 (2004).